# Digital Quantum Simulation of the Nonlinear Lindblad Master Equation Based on Quantum Trajectory Averaging

### v1

@@ Abstract
Since precisely controlling dissipation in realistic environments is challenging, digital simulation of the Lindblad master equation (LME) is of great significance for understanding nonequilibrium dynamics in open quantum systems. However, achieving long-time simulations for complex systems with multiple dissipation channels remains a major challenge, both theoretically and experimentally. Here, we propose a 2-dilation digital simulation scheme for the non-linear Lindblad master equation (NLME) based on quantum trajectory averaging. The NLME continuously interpolates between full LME and the dynamical equation governed by the effective non-Hermitian Hamiltonian. Remarkably, for standard LMEs, our scheme reduces to a 1-dilation method that enables deterministic realizations without postselection. This deterministic nature overcomes a key limitation in some existing simulation methods, where repeated postselections lead to exponentially vanishing implementation probabilities. Consequently, our scheme allows efficient long-time simulations of LMEs with multiple jump operators. As a demonstration, we present numerical experiments simulating novel theoretical predictions in open quantum systems, including localization in open quantum systems and the postselected skin effect.
@@
Introduction.— The Lindblad master equation (LME) plays a crucial role in the study of open quantum systems, providing a rigorous framework to describe the quantum system coupling with the environment under the Born-Markov approximation Lindblad1976 ; Gardiner1985 ; Daniel2020 . In recent years, the demand for precise control over dissipative processes has driven extensive theoretical efforts to explore novel nonequilibrium dynamical phenomena within the Lindblad framework, such as Liouvillian flat bands Hugo2023 , postselected skin effects Hugo2025 , and localization in open quantum systems Yusipov2017 ; YLiu2024 . These studies often require carefully engineered dissipation operators, which are difficult to implement on analog simulation platforms. In contrast, digital quantum platforms provide a standardized and promising approach to simulation.
@@
The primary challenge of digitally simulating the LME lies in implementing its intrinsically non-unitary nature using unitary logic gates. This difficulty can be addressed through approaches such as dilation methods RCleve2017 ; ZDing2024 ; ZHu2020 ; KMarsden2021 ; ZHu2022 ; ASchlimgen2021 ; ASchlimgen2022 ; NSuri2023 ; DMazziotti2022 ; EOh2024 ; SPeng2024 ; LHu2019 ; JHan2021 , which effectively embed non-unitarity into enlarged unitary gates, and hybrid quantum-classical algorithms, including variational quantum simulation XYuan2019 ; SEndo2020 ; Mahdian2020 ; JLuo2024 and the quantum imaginary time evolution method MMotta2020 ; HKamakari2022 . Currently, effective experimental schemes have been demonstrated for single-qubit and few-qubit systems LHu2019 ; JHan2021 ; ASchlimgen2021 ; ZHu2020 ; ZHu2022 ; EOh2024 ; HKamakari2022 ; JWPan2017 . However, for large systems, numerous dissipation sources, and long-time evolution, significant theoretical and experimental challenges remain. For instance, the Stinespring dilation approach RCleve2017 ; ZDing2024 requires a large number of auxiliary qubits, potentially affecting its scalability. Moreover, if probabilistic approaches, such as Sz.-Nagy dilation (SND) ZHu2020 ; KMarsden2021 ; ZHu2022 , the linear combination of unitaries (LCU) method ASchlimgen2021 ; ASchlimgen2022 ; NSuri2023 , and singular-value decomposition (SVD) method DMazziotti2022 ; EOh2024 are used to iteratively simulate system evolution within each time step $\delta t$, the overall success probability rapidly approaches zero as the number of time steps and dissipation sources increases.
An alternative simulation scheme for the LME on a classical computer is the quantum trajectory method Dalibard1992 ; Dum1992 ; Gardiner1992 ; Klaus1993 ; Plenio1998 ; Daley2014 ; HWiseman2016 , which interprets the LME as a stochastic average over individual trajectories of pure states. This approach significantly reduces memory usage, as it focuses on the state rather than the density matrix. Inspired by this method, Luo et al. JLuo2024 and Peng et al. SPeng2024 recently proposed new variational and dilation-based simulation schemes, which substantially reduce logical gate complexity. However, the former requires additional ansatz assumptions, while the latter remains a probabilistic implementation and still faces the challenge of an almost vanishing success probability at long times.
@@
In our trajectory simulation scheme, we overcome the challenge of the overall success probability approaching zero. The key insight is to rigorously match the stochasticity in quantum trajectories with the probabilistic outcomes of measurements in quantum circuits. Each measurement outcome directly corresponds to a specific mathematical process in the quantum trajectory method: either a quantum jump or non-unitary evolution. Unlike methods based on SND, LCU, and SVD, where certain states are discarded after measurement, our approach ensures that every measurement outcome is meaningful.
Based on the principle of efficiently utilizing measurement randomness, our scheme can extend from deterministic trajectory simulations of the LME to probabilistic trajectory simulations of the non-linear Lindblad master equation (NLME) Hugo2025 ; Lewalle2023 ; MCoppola2024 ; APaviglianiti2024 . The NLME describes the dynamics of an open system lacking an arbitrary proportion of quantum jumps, providing a continuous interpolation between the LME and the evolution of effective non-Hermitian Hamiltonian (ENHH). It gives a reasonable framework for studying the interplay and competition between the non-Hermitian Hamiltonians and pure dissipation or pure measurement processes.
In this paper, we first introduce the trajectory-averaged interpretation of the NLME. Based on this interpretation, we propose a probabilistic 2-dilation implementation. For two special cases of the NLME, namely the LME and the ENHH, the scheme reduces to a 1-dilation. In particular, for the LME, the scheme becomes fully deterministic due to the absence of post-selection effects. Furthermore, we simulate a specific class of jump operators that are theoretically interesting but difficult to realize experimentally, demonstrating a numerical approach for simulating open-system localization and the postselected skin effect.
@@
Figure 1: Scheme of the quantum trajectory simulation: (a) the 2-dilation method for the NLME as Eq. (1), (b) the 1-dilation method for the LME as Eq. (2), and (c) the 1-dilation method for the evolution of ENHH as Eq. (3). The system is initialized in state $|\phi(0)\rangle$. The auxiliary qubits are initialized in state $|00\rangle_{a}$ in (a) and in state $|0\rangle_{a}$ in both (b) and (c). At each time step $\delta t$, the operations inside the brackets are executed. The gate $U_{0}=e^{-iH\delta t}$ represents the evolution from system Hamiltonian. The gate $U_{\mu\geq 1}$, corresponding to the $\mu$-th dissipation source, is given by Eq. (7) in (a), Eq. (11) in (b), and Eq. (12) in (c). After each measurement of auxiliary qubit, the postselection, denoted by ’P’, is required only in (a) and (c), making their realization probabilistic. In contrast, the method in (b) is deterministic as it does not require postselection. In all three cases, the auxiliary qubit must be repeatedly reset to its initial state, denoted by ’R’.
@@
The NLME can be interpreted by the quantum trajectory method Hugo2025 . The expectation value of an arbitrary observable $\hat{O}$ is given by the average over many trajectories: $\langle\hat{O}\rangle=\mathrm{Tr}(\hat{O}\rho)=\lim_{K\to\infty}\frac{1}{K}\sum_{i=1}^{K}\langle\phi_{i}(t)|\,\hat{O}\,|\phi_{i}(t)\rangle$. The quantum state at time $t+\delta t$ for the $i$-th trajectory is obtained from the state at time $t$ through a stochastic evolution:
$\begin{split}&|\phi_{i}(t+\delta t)\rangle=e^{-iH\delta t}\prod_{\mu=1}^{M}\mathcal{U}_{\mu}(\delta t)\ |\phi_{i}(t)\rangle,\\ &\mathcal{U}_{\mu}(\delta t)=\left\{\begin{matrix}(1-\eta_{\mu})\delta p_{\mu}:&\mathcal{N}L_{\mu}\\ 1-(1-\eta_{\mu})\delta p_{\mu}:&\mathcal{N}\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\end{matrix}\right.\ ,\end{split}$ (4)
@@
where the subscript $a$ represents auxiliary qubits. After measuring the auxiliary qubits, the system probabilistically collapses into the state of $\mathcal{N}X_{\mu}|\phi\rangle$ ($X$ = $A$, $B$, or $C$) with a probability of $\langle X_{\mu}^{\dagger}X_{\mu}\rangle:=\langle\phi|X_{\mu}^{\dagger}X_{\mu}|\phi\rangle$.
@@
$\begin{split}&A_{\mu}=\sqrt{1-\gamma_{\mu}\delta tL_{\mu}^{\dagger}L_{\mu}}\,,\ \ \tilde{A}_{\mu}=\sqrt{1-\gamma_{\mu}\delta tL_{\mu}L_{\mu}^{\dagger}}\,,\\ &B_{\mu}=\sqrt{(1-\eta_{\mu})\gamma_{\mu}\delta t}\,L_{\mu}\,,\ \ C_{\mu}=\sqrt{\eta_{\mu}\gamma_{\mu}\delta t}\,L_{\mu}\,,\end{split}$ (8)
@@
The 2-dilation method for realizing the full trajectory is presented in Fig. 1(a). The system and auxiliary qubits are initialized in the state $|\phi_{i}(0)\rangle$ and $|00\rangle_{a}$. Then, N iterations are carried out for a time evolution of duration $T=N\delta t$. During each time step $\delta t$, the unitary evolution governed by the system Hamiltonian is encoded in the logic gate $U_{0}=e^{-iH\delta t}$. Each dissipation effect $\mathcal{U}_{\mu}(\delta t)$ is implemented by measuring and post-selecting the auxiliary qubits after applying the 2-dilation logic gate $U_{\mu}$ in Eq. (7), where only the measurement outcomes of $|01\rangle_{a}$ and $|10\rangle_{a}$ are considered valid. If the outcome is $|00\rangle_{a}$ the iteration is ended, and this trajectory is discarded. Following postselection, the auxiliary qubits are reset to $|00\rangle_{a}$ state in preparation for the next gate $U_{\mu+1}$.
@@
$P$ is roughly estimated as $(1-\bar{\eta}\bar{\gamma}\langle L_{\mu}^{\dagger}L_{\mu}\rangle\delta t)^{NM}$, where $\bar{\eta}=\frac{1}{M}\sum_{\mu=1}^{M}\eta_{\mu}$ and $\bar{\gamma}=\frac{1}{M}\sum_{\mu=1}^{M}\gamma_{\mu}$.
@@
Numerical experiments.— We first simulate a two-level atom with monitoring its spontaneous emission Hugo2025 ; SuMa to demonstrate the validity of our scheme. The dynamic equation is
${}\begin{split}\frac{d}{dt}\rho=&-iJ[\sigma^{x},\rho]\\ &+\gamma\left(-\frac{1}{2}\{\sigma^{+}\sigma^{-},\rho\}+(1-\eta)\sigma^{-}\rho\sigma^{+}+\eta\langle\sigma^{+}\sigma^{-}\rangle\rho\right),\end{split}$ (13)
where $\sigma^{\pm}$ is the spin raising and lowering operators and $\sigma^{x}=\sigma^{+}+\sigma^{-}$. This NLME can be exactly solved by the vectorization method Hugo2025 ; SuMa . We compare the evolution of excited-state probability from the exact solution with that from the digital simulation of our 2-dilation and 1-dilation methods in Fig. 1, where the curves exhibit good consistency.
[…]
The fundamental mission of quantum simulation is to explore complex systems. Now, we focus on a many-particle problem in an L-site open chain, with the evolution equation given by:
$\begin{split}~{}&\frac{d}{dt}\rho=-i[H,\rho]+\gamma\sum_{l=1}^{L-1}\mathcal{D}_{l}(\rho),\\ &H=J\sum_{l=1}^{L-1}\left(\sigma_{l}^{+}\sigma_{l+1}^{-}+{\rm H.c.}\right)+\sum_{l=1}^{L}V\cos(2\pi\omega l)\sigma_{l}^{z},\\ &\mathcal{D}_{l}(\rho)=-\frac{1}{2}\{L_{l}^{{\dagger}}L_{l},\rho\}+(1-\eta)L_{l}\rho L_{l}^{{\dagger}}+\eta\langle L_{l}^{{\dagger}}L_{l}\rangle\rho,\end{split}$ (14)
where $\sigma_{l}^{\mu}(\mu=z,+,-)$ is the spin-1/2 operator at the site $l$, and $\omega=(\sqrt{5}-1)/2$ indicates the quasicrystal potential. The local dissipation operator takes the following form:
${}L_{l}=\frac{1}{2}(\sigma_{l}^{+}+e^{i\alpha}\sigma_{l+1}^{+})(\sigma_{l}^{-}+e^{i\beta}\sigma_{l+1}^{-}),$ (15)
which represents the phase-changing effect between neighboring spins. This kind of dissipation operator exhibits novel phenomena in theoretical studies, such as entangled state preparation BKraus2008 ; SDiel2008 ; PSchindler2013 , localization Yusipov2017 ; YLiu2024 , and the skin effect in open quantum systems Hugo2025 ; Wang2022 ; Feng2023 ; ZCLiu2024 . Note that this model is an intrinsic many-body evolution due to the $\hat{n}_{l}\,\hat{n}_{l+1}$ interaction ($\hat{n}_{l}:=\sigma_{l}^{+}\sigma_{l}^{-}$) existing in the evolution factor of $-\frac{1}{2}\gamma L_{l}^{\dagger}L_{l}\delta t$.
[…]
Furthermore, when $\alpha=\pi/2$, $\beta=-\pi/2$, $V=0$, and $\eta\neq 0$, the steady state is expected to exhibit skin effect Hugo2025 . We simulate this postselected skin effect by our 2-dilation method in the supplementary materials SuMa .

### v2

# Digital Quantum Simulation of the Lindblad Master Equation and Its Nonlinear Extensions via Quantum Trajectory Averaging

@@ Abstract
Since precisely controlling dissipation in realistic environments is challenging, digital simulation of the Lindblad master equation (LME) is of great significance for understanding nonequilibrium dynamics in open quantum systems. However, achieving long-time simulations for complex systems with multiple dissipation channels remains a major challenge, both theoretically and experimentally. Here, we propose a 1-dilation digital scheme for simulating the LME based on quantum trajectory averaging without postselection. By rigorously matching the stochasticity inherent in quantum trajectories with the probabilistic outcomes of quantum measurements, our method effectively translates the classically established quantum jump algorithm into executable quantum circuits. A key advantage of our method is that it overcomes the exponential suppression of success probability seen in some existing postselection-dependent schemes, especially for long-time evolution or systems with numerous jump operators. Moreover, the scheme can be extended to a 2-dilation framework for the nonlinear LME with postselection, bridging the full LME and non-Hermitian Hamiltonian dynamics. This extended scheme provides a digital approach for exploring the interplay between non-Hermitian Hamiltonians and dissipative terms within a monitored quantum dynamics framework.
@@
Introduction.— The Lindblad master equation (LME) plays a crucial role in the study of open quantum systems, providing a rigorous framework to describe the quantum system coupling with the environment under the Born-Markov approximation Lindblad1976 ; Gardiner1985 ; Daniel2020 . In recent years, the demand for precise control over dissipative processes has driven extensive theoretical efforts to explore novel nonequilibrium dynamical phenomena within the Lindblad framework, such as Liouvillian flat bands Hugo2023 , postselected skin effects Hugo2025 , and localization in open quantum systems Yusipov2017 ; Vakulchyk2018 ; YLiu2024 . These studies often require carefully engineered dissipation operators, which are difficult to implement on analog simulation platforms. In contrast, digital quantum platforms provide a standardized and promising approach to simulation.
@@
The primary challenge of digitally simulating the LME lies in implementing its intrinsically non-unitary nature using unitary logic gates. This difficulty can be addressed through approaches such as dilation methods RCleve2017 ; ZDing2024 ; ZHu2020 ; KMarsden2021 ; ZHu2022 ; ASchlimgen2021 ; ASchlimgen2022 ; NSuri2023 ; DMazziotti2022 ; EOh2024 ; SPeng2025 ; LHu2019 ; JHan2021 ; WJYu2025 , which effectively embed non-unitarity into enlarged unitary gates, and hybrid quantum-classical algorithms, including variational quantum simulation XYuan2019 ; SEndo2020 ; Mahdian2020 ; JLuo2024 and the quantum imaginary time evolution method MMotta2020 ; HKamakari2022 . However, for large systems, numerous dissipation sources, and long-time evolution, significant theoretical and experimental challenges remain. For instance, probabilistic simulation approaches relying on postselection, such as those based on Sz.-Nagy dilation (SND) ZHu2020 ; KMarsden2021 ; ZHu2022 , the linear combination of unitaries (LCU) method ASchlimgen2021 ; ASchlimgen2022 ; NSuri2023 , and singular-value decomposition (SVD) DMazziotti2022 ; EOh2024 , face the challenge of a rapidly diminishing overall success probability as the number of time steps or dissipation sources increases. For methods that avoid postselection, such as the one proposed by Ding et al. ZDing2024 based on the Stinespring dilation of Kraus operators, the critical issue often lies in the frequent requirement for larger number of auxiliary qubits, which can potentially affect the scalability of the simulation, especially on near-term quantum hardware with limited qubit counts. Recently, Peng et al. SPeng2025 proposed a trajectory-inspired Lindbladian simulation that reduces the circuit complexity with respect to the number of jump operators. However, it still relies on oblivious amplitude amplification to overcome the postselection problem, which increases the circuit depth and the required control precision.
@@
Here, we propose a postselection-free Lindbladian simulation method. It avoids the challenge of vanishing success probability using only one ancilla qubit. Our approach provides a translation of the classically known quantum jump algorithm Dalibard1992 ; Dum1992 ; Klaus1993 ; Plenio1998 ; Daley2014 into executable quantum circuits. The key insight, as shown in Fig. 1, is to exactly match the stochasticity in quantum trajectories with the probabilistic outcomes of measurements in quantum circuits: each measurement result directly maps to a specific mathematical process in the quantum trajectory method, representing either a quantum jump or a non-unitary evolution step. In contrast to methods such as SND, LCU, and SVD, where certain measurement outcomes must be discarded, our protocol ensures that every measurement outcome is meaningful and contributes to the simulation.
Moreover, our scheme can be extended to probabilistically simulate a class of nonlinear Lindblad master equations (NLME) Hugo2025 ; Lewalle2023 ; MCoppola2024 ; APaviglianiti2025 using two ancilla qubits. The NLME continuously interpolates between the LME and the evolution equation of effective non-Hermitian Hamiltonian (ENHH). It arises in monitored open quantum systems where quantum jumps are lost or postselected. Such equations provide a theoretical framework for probing the interplay between non-Hermitian effects and dissipation Hugo2025 , as well as measurement-induced phenomena in imperfect settings MCoppola2024 ; APaviglianiti2025 , processes of information loss APaviglianiti2025 and quantum optimal control Lewalle2023 . Conventional experimental proposals for realizing NLMEs rely on imperfect measurement processes Hugo2025 ; Minganti2020 , which often encounter practical limitations due to uncontrollable efficiency of monitors. In contrast, our digital simulation scheme provides a flexible and highly controllable alternative.
Figure 1: Realizing stochastic quantum trajectories via quantum circuit measurements. (a) In the classical algorithm, the system undergoes Hamiltonian evolution (black) followed by a stochastic choice between non-unitary evolution (red double lines) and quantum jumps $L_{\mu}$ (blue). The branching probabilities are determined by Eq. (4). (b) In the circuit implementation, the unitary $U_{0}$ corresponds to the Hamiltonian part, while the operators $A$ and $B$ (Eq. (8)) represent the non-unitary and jump components. Ancilla measurements (black triangles) facilitate the stochasticity for LME, while NLME implementation further requires postselection on the measurement outcomes.
In this paper, we first present a trajectory-averaged interpretation of the NLME and propose a probabilistic 2-dilation implementation, which reduces to a 1-dilation for special cases (LME and ENHH). For LME, the scheme becomes fully deterministic due to absent post-selection effects. We analyze error scaling and validate the approach through dissipative XXZ model simulations. We further perform exploratory simulations of novel models featuring many-body localization in open systems and the postselected skin effect to demonstrate quantum computing’s potential in investigating emerging phenomena in the Supplementary Material SuMa .
@@
Figure 2: Scheme of the digital quantum trajectory simulation: (a) the 2-dilation method for the NLME as Eq. (1), (b) the 1-dilation method for the LME as Eq. (2), and (c) the 1-dilation method for the evolution of ENHH as Eq. (3). The system is initialized in state $|\phi(0)\rangle$. The auxiliary qubits are initialized in state $|00\rangle_{a}$ in (a) and in state $|0\rangle_{a}$ in both (b) and (c). At each time step $\delta t$, the operations inside the brackets are executed. The $U_{0}$ represents the Hamiltonian simulation of the time evolution $e^{-iH\delta t}$. The gate $U_{\mu\geq 1}$, corresponding to the $\mu$-th dissipation source, is given by Eq. (7) in (a), Eq. (11) in (b), and Eq. (12) in (c). After each measurement of auxiliary qubit, the postselection, denoted by ’P’, is required only in (a) and (c), making their realization probabilistic. In contrast, the method in (b) is deterministic as it does not require postselection. In all three cases, the auxiliary qubit must be repeatedly reset to its initial state, denoted by ’R’.
@@
The NLME can be interpreted by the quantum trajectory method Hugo2025 . The expectation value of an arbitrary observable $\hat{O}$ is given by the average over many trajectories: $\langle\hat{O}\rangle=\mathrm{Tr}(\hat{O}\rho)=\lim_{K\to\infty}\frac{1}{K}\sum_{i=1}^{K}\langle\phi_{i}(t)|\,\hat{O}\,|\phi_{i}(t)\rangle$. The quantum state at time $t+\delta t$ for the $i$-th trajectory is obtained from the state at time $t$ through a stochastic evolution, as shown in Fig. 1(a):
[…]
@@
where the subscript $a$ represents auxiliary qubits, and the state $|11\rangle_{a}$ is forbidden by the structure of $U_{\mu}$. After measuring the auxiliary qubits, the system probabilistically collapses into the state of $\mathcal{N}X_{\mu}|\phi\rangle$ ($X$ = $A$, $B$, or $C$) with a probability of $\langle X_{\mu}^{\dagger}X_{\mu}\rangle:=\langle\phi|X_{\mu}^{\dagger}X_{\mu}|\phi\rangle$.
@@
[…]
@@
The 2-dilation method for realizing the full trajectory is presented in Fig. 2(a). The system and auxiliary qubits are initialized in the state $|\phi_{i}(0)\rangle$ and $|00\rangle_{a}$. Then, $N$ iterations are carried out for a time evolution of duration $T=N\delta t$. During each time step $\delta t$, the unitary evolution $e^{-iH\delta t}$ governed by the system Hamiltonian $H$ is encoded in the logic gate $U_{0}$, which is an implemention by the Hamiltonian simulation Childs2018 (e.g., via Trotterization). Each dissipation effect $\mathcal{U}_{\mu}(\delta t)$ is implemented by measuring and post-selecting the auxiliary qubits after applying the 2-dilation logic gate $U_{\mu}$ in Eq. (7), where only the measurement outcomes of $|01\rangle_{a}$ and $|10\rangle_{a}$ are considered valid. If the outcome is $|00\rangle_{a}$ the iteration is ended, and this trajectory is discarded. Following postselection, the auxiliary qubits are reset to $|00\rangle_{a}$ state in preparation for the next gate $U_{\mu+1}$.
@@
This implies that the success probability follows an approximately exponential decay as a function of $T\times M\times\eta$.
@@
Error analysis and algorithmic scaling.— The errors in our quantum simulation schemes primarily stem from two sources: the random sampling error introduced during trajectory averaging, and the systematic error arising from the non-commutativity of operators within the $\delta t$ interval. According to the central limit theorem, the sampling error decreases as $1/\sqrt{K}$, where $K$ is the number of trajectories. Therefore, for bounded observables, the sampling error can be suppressed below the level of the systematic error by increasing the number of trajectories.
For the systematic error, our algorithm maintains first-order accuracy within each $\delta t$. Consequently, for an NLME simulation of total time $t$ targeting a precision $\varepsilon$, the required circuit depth scales as $\mathcal{O}(t^{2}/\varepsilon)$. For the LME simulation, further results can be obtained through diamond norm analysis (see Supplementary Material SuMa ). Specifically, when $U_{0}$ is first-order accurate in $\delta t$, the scaling behavior of the circuit depth is
$\mathcal{O}\left(M\frac{\lambda^{2}t^{2}}{\varepsilon}\right).$ (13)
Here, $M$ is the number of dissipative channels, and $\lambda=\lVert H\rVert+M\gamma_{x}\lVert L_{x}\rVert^{2}$, where $\gamma_{x}\lVert L_{x}\rVert^{2}$ is the maximum value in the set of spectral norms: $\{\gamma_{1}\lVert L_{1}\rVert^{2},\gamma_{2}\lVert L_{2}\rVert^{2},\cdots\gamma_{M}\lVert L_{M}\rVert^{2}\}$.
Numerical experiments.— We numerically test our scheme by simulating different dynamical processes. A detailed pedagogical example of a two-level atom monitored through its spontaneous emission Hugo2025 is provided in the Supplemental Material SuMa , to demonstrate the validity of our NLME simlation approach.
Here, we present results for a dissipative XXZ spin chain governed by the LME:
$\frac{d}{dt}\rho=-i[H,\rho]+\gamma\sum_{l=1}^{L}\left(-\frac{1}{2}\{\sigma_{l}^{+}\sigma_{l}^{-},\rho\}+\sigma_{l}^{-}\rho\sigma_{l}^{+}\right),$ (14)
with the Hamiltonian
$H=J\sum_{l=1}^{L-1}\left(\sigma_{l+1}^{x}\sigma_{l}^{x}+\sigma_{l+1}^{y}\sigma_{l}^{y}+\Delta\sigma_{l+1}^{z}\sigma_{l}^{z}\right),$ (15)
where $\sigma_{l}^{\pm}$ and $\sigma_{l}^{x,y,z}$ are the spin raising/lowering operators and Pauli matrices at site $l$, respectively. We compare the numerical simulation of the Eq. (14) presented in Fig. 2(b) with the exact solutions, as shown in Fig. 3, where the $U_{0}$ is implemented using a second-order Trotter-Suzuki decomposition for the Hamiltonian in Eq. (15) split into even and odd terms. As shown in Fig. 3(a), the numerical results for the spin-up probability at the first site ($\rm P_{1}$) and the averaged nearest-neighbor correlations ($C_{zz}=\frac{1}{L-1}\sum_{j=1}^{L-1}\langle\sigma_{l+1}^{z}\sigma_{l}^{z}\rangle$) agree well with the exact solutions. The associated standard error decreases significantly as time and the number of trajectories increase. Meanwhile, we study the dependence of the numerical error of the density matrix on the time step $\delta t$. Here, the error is defined as the trace norm of the difference between $\rho^{\prime}$ obtained from our trajectory-based method and the exact $\rho$, i.e., $error=\lVert\rho^{\prime}-\rho\rVert_{1}$, evaluated at the final time $T=10$. As shown in Fig. 3(b), the error decreases as $\delta t$ becomes smaller, nearly scaling as $\delta t^{0.8}$.
Figure 3: Numerical simulation of dissipative XXZ spin chain. We study the evolution of a 5-site dissipative XXZ chain from $t=1$ to 1010. The system is initialized in the all-spin-up state, with parameters $(J,\Delta,\gamma)=(1,2,0.5)$. (a) Exact results of the spin-up probability at the first site, $P_{1}={\rm Tr}(\sigma_{1}^{+}\sigma_{1}^{-}\rho)$, and the averaged nearest-neighbor correlations, $C_{zz}=\frac{1}{L-1}\sum_{j=1}^{L-1}\langle\sigma_{l+1}^{z}\sigma_{l}^{z}\rangle$, are shown as the blue dashed line and the solid magenta line, respectively. Circular and rhombus markers with error bars denote the results of our digital simulation method with time step $\delta t=0.1$, using 100 and 1000 trajectories, respectively. (b) Scaling of the error of the density matrix $\rho$ at $t=10$ with the simulation time step $\delta t$, obtained by averaging over 100,000 trajectories. Data ponts correspond to $\delta t=(1.0,0.5,0.2,0.1,0.05,0.02,0.01)$, and the solid line shows a fit to the data.
Furthermore, we explore two complicated phenomena: the postselected skin effect Hugo2025 and open-system localization Yusipov2017 ; Vakulchyk2018 ; YLiu2024 in the Supplementary Material SuMa . These remain challenging to address with current analog simulations or classical computation. The latter, in particular, involving dissipation-induced interactions, could be considered as an open question of many-body localization in open system. The preliminary small-scale simulation results are provided to imply the potential advantage of quantum computing in exploring novel problems.